40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 24 00:39:47 2017, gautam, Update, LSC, DRMI locking attempt 
    Reply  Fri Aug 25 01:20:52 2017, gautam, Update, LSC, DRMI locking attempt 
       Reply  Wed Aug 30 19:49:33 2017, gautam, Update, LSC, REFL55 demod board debugging REFL55_demod_check.pdf
          Reply  Thu Aug 31 00:52:52 2017, gautam, Update, LSC, REFL55 whitening board debugging REFL55_whtCheck.pdfREFL55_whtChk.tar.gz
             Reply  Thu Aug 31 03:31:15 2017, gautam, Update, LSC, DRMI re-locked! DRMI_1f.pngDRMI_relocked.pdf
                Reply  Mon Sep 4 16:30:06 2017, gautam, Update, LSC, Oplev loop tweaking 
                   Reply  Mon Sep 4 18:18:29 2017, rana, Update, LSC, dewhite switching: FOTON settings 
                   Reply  Tue Sep 5 02:07:49 2017, gautam, Update, LSC, Low Noise DRMI attempt MICH_err_comp.pdfdeWhitenedCoil.pdf
                      Reply  Wed Sep 6 23:06:30 2017, gautam, Update, LSC, Coil de-whitening switching investigation BScoils.pdfITMcoils.pdf
                         Reply  Fri Sep 15 16:00:33 2017, gautam, Update, LSC, Sensing measurement DRMI1f_Sep5.pdf
                         Reply  Fri Sep 15 17:08:58 2017, gautam, Update, LSC, Coil de-whitening switching investigation DRMI_IntensityNoise.pdf
                            Reply  Sat Sep 16 10:56:19 2017, rana, Update, LSC, Coil de-whitening switching investigation 
                               Reply  Fri Sep 22 18:12:27 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                  Reply  Wed Sep 27 22:25:21 2017, gautam, Update, LSC, DAC noise measurement (again) coilNoises.pdf
                                     Reply  Thu Sep 28 06:35:44 2017, Chris, Update, LSC, DAC noise measurement (again) 
                                        Reply  Thu Sep 28 23:50:04 2017, gautam, Update, LSC, DAC noise measurement (again) 
                                           Reply  Fri Sep 29 18:36:25 2017, gautam, Update, LSC, DAC noise measurement (again) 
                               Reply  Tue Oct 3 01:32:39 2017, gautam, Update, LSC, Laser intensity noise coupling to MICH (simulated) MICH_intensityNoiseCoupling.pdfMICH_intensityNoiseASD.pdf
Message ID: 13248     Entry time: Thu Aug 24 00:39:47 2017     Reply to this: 13252
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking attempt 

Since the single arm locking and dither alignment seemed to work alright after the CDS overhaul, I decided to try some recycling cavity locking tonight.

  • First, I locked single arms, ran dither alignment servos, and centered all test mass Oplevs. Note: the X arm dither alignment doesn't seem to work if we use the High-Gain Thorlabs PD as the Transmission PD. The BS loops just seem to pick up large offsets and the alignment actually degrades over a couple of minutes. This needs to be investigated.
  • Next, to get good PRM alignment, I manually moved the EPICS sliders till the REFL spot became roughly centered on the CCD screen.
  • Then I tried locking PRMI on carrier using the usual C1IFOConfigure script - the lock was caught within ~30 seconds.
  • The PRCL and MICH dither servo scripts also ran fine.
    • Centered PRM Oplev.
  • Next, I tried enabling the PRC angular feedforward.
    • OAF model does not automatically revert to its safe.snap configuration on model reboot, so I first manually did this such that the correct filter banks were enabled.
    • I was able to turn on the angular feedforward without disturbing the PRMI carrier lock. The angular motion of the POP spot on the CCD monitor was visibly reduced.
  • At this point I decided to try DRMI locking.
    • I centered the beam on the AS PDs with the simple Michelson.
    • Centered the beam on the REFL PDs with PRM aligned and PRC flashing through resonances.
    • Restored SRM alignment by eye with EPICS sliders.
    • Cavity alignment seemed alright - so I tried to lock DRMI with the old settings (i.e. from DRMI 1f locking a couple of months ago). But I had no success.
    • The behaviour of REFL55 (used for SRCL control) has changed dramatically - the analog whitening gain for this PD used to be +18dB, but at this setting, there are frequent ADC overflows. I had to reduce the whitening gain to +6dB to stop the ADC overflows. I also checked to make sure that the whitening setting was "manual" and not triggered.

Why should this have changed? I was just on the AS table and did re-center the beam onto the REFL 55 RFPD, but I had also done this in April/May when I was last doing DRMI locking. But I can't explain the apparent factor of ~4 increase in light level. I think I have some measurements of the light levels at various PDs from April 2017, I will see how the present levels line up.

Of course dataviever won't cooperate when I am trying to monitor testpoints.

I may be missing something obvious, but I am quitting for tonight, will look into this more tomorrow.


Unrelated to this work: looking at the GTRY spot on the CCD monitor, there seems to be some excess angular motion. Not sure where this is coming from. In the past, this sort of problem has been symptomatic of something going wonky with the Oplev loops. But I took loop measurements for ITMY and ETMY PIT and YAW, they look normal. I will investigate further when I am doing some more ALS work.

ELOG V3.1.3-